OpenStudy (anonymous):

(-47x^2y/141x^3y^-3 simplify into a positive exponent

OpenStudy (anonymous):

-47/141 = -1/3 x^2/x^3 = x^-1 y/y^-3 = y^4 so -1 y^4/ 3x

hero (hero):

\[\frac{-47x^2y^4}{141x^3} \]

hero (hero):

And the simplified version of that is... \[-\frac{y^4}{3x}\]